# MG-SLAM: Structure Gaussian SLAM with Manhattan World Hypothesis

### v2

@@ 4.3 Evaluation on large-scale dataset
To evaluate the robustness of the system in large-scale indoor environments, we evaluate MG-SLAM on the Replica Apartment dataset [38]. This dataset contains extensive multi-room scenes, complex object geometries, and looping trajectories across rooms. Table 3 presents the rendering quality of our method compared to SplaTAM [19] and MonoGS [28] over five selected scenes. MG-SLAM shows a notable improvement over these baselines, particularly achieving a 7dB improvement in the extensive apartment 0 scene, which features a two-floor, eight-room layout. This optimal performance is largely attributed to the inclusion of the fused line segments, which lays a solid foundation for loop closure and pose optimization. Figure 5 illustrates the results of novel-view rendering for apartment scenes. Our method shows remarkable enhancements over Gaussian baselines in geometry accuracy and the richness of fine details.
4.4 Evaluation of scene completion
The Gaussian SLAM faces limitations in inferring the geometry of unseen views. This issue is especially evident in indoor scenes that feature complex geometry, where basic surfaces like floors are often obscured and poorly represented. Figure 6 compared our method with SplaTAM [19] and MonoGS [28] in scenes with missing geometry. Utilizing the surface optimization strategy based on the Manhattan World hypothesis [5], our method can accurately detect and proactively generate new Gaussians efficiently to fill gaps with certain textures, whereas Gaussian baseline methods exhibit substantial defects on structured surfaces. Demonstrations for real-world scenes can be found in Section D.4.
Figure 6: Qualitative comparison of our method and the Gaussian SLAM baselines for reconstruction across selected scenes from the Replica dataset [38].
5 Conclusion
In this study, we present MG-SLAM, a Gaussian-based SLAM method based on the Manhattan World hypothesis. MG-SLAM employs line segments for robust pose estimation and map refinement. Furthermore, by leveraging the line segments and planar assumption, we efficiently interpolate new Gaussians on gaps of missing geometry. Extensive experiments demonstrate that our method delivers state-of-the-art tracking and mapping performance, while also achieving remarkable high speed detailed in Section D.5. Limitations of our work are discussed in Appendix E.
Appendix A Implementation details
In this section, we describe the implementation details of our systems. Our tracking system is implemented in C++, while the mapping system uses Python3 and CUDA C. The experiments were conducted using a single RTX A100-80GB GPU and 24-core AMD EPYC 7402 processor. We use Adam optimizer for Gaussian representation optimization and network training. The hyperparameters for 3D Gaussian splitting [20] are the same as the original paper. Table 4 presents the values for hyperparameters in our system. For Gaussian insertion, in each keyframe, we sampled $\mu_{\mathcal{G}}$ in Equation 8 from the rendered depth map $\mathcal{D}_{\rm pix}$ following the distribution of $\mathcal{N}(\mathcal{D}_{\rm pix},0.2\sigma_{\mathcal{D}})$. For uninitialized areas, we initialize Gaussians by sampling from $\mathcal{N}(\bar{\mathcal{D}}_{\rm pix},0.5\sigma_{\mathcal{D}})$ where $\bar{\mathcal{D}}_{\rm pix}$ is the mean value. For pruning, we remove the Gaussians with opacity less than 0.6. Our system incorporates a segmentation loss introduced in [26], and we leverage this semantic information to effectively extract structured surfaces, such as floors and ceilings, from the reconstructed scenes. This segmentation sufficiently fulfills our objective to extract the current surfaces and train a PointNet++ [34] network to predict the textures of interpolated Gaussians, which are used to fill the missing geometry.
Table 4: MG-SLAM hyperparameters
Symbol Explanation Value
$N_{p}$ number of point features in tracking 4e3
$N_{\ell}$ number of line segments in tracking 200
$r_{d}$ down-sample ratio for Gaussian sampling from point cloud 8
$n_{map}$ mapping iteration 150
$SH_{degree}$ degree of the spherical harmonics 0
$l_{feat}$ learning rate for SH features 2.5e-3
$l_{opacity}$ learning rate for opacity 5e-2
$l_{scale}$ learning rate for scaling 1e-3
$l_{rotat}$ learning rate for rotation 1e-3
$\lambda_{\mathcal{D}}$ weighting term for depth loss in mapping 0.1
$\lambda_{\mathcal{S}}$ weighting term for semantic loss in mapping 0.1
$\lambda_{\ell}$ re-weighting term for line feature loss 0.25
$\mathcal{T}_{\ell}$ line segment filtering threshold (relative to image scale) 8e-2
$\mathcal{T}_{density}$ density threshold for hole detection 0.9
$b_{\mathcal{G}}$ batch size of Gaussian partition to train PointNet++ [34] 1e4
$n_{e}$ number of epoch to train PointNet++ [34] 25
Appendix B Additional methods
B.1 Gaussian density
In our Gaussian scenes, we define the density of the Gaussian $\nu:\mathbb{R}^{3}\rightarrow\mathbb{R}_{+}$ as the sum of the Gaussian values weighted by their alpha-blending coefficients at any given grid points p as:
where the $\mu_{\mathcal{G}_{i}},\Sigma_{\mathcal{G}_{i}},\alpha_{\mathcal{G}_{i}}$ are the means, covariances, and alpha-blending coefficients of the Gaussians. To simplify the calculation, $\nu(p)$ can be approximated by:
where $\mathcal{G}^{*}$ is the nearest Gaussian that contributes most to $p$. This Gaussian density function $\nu^{*}(\cdot)$ facilitates the identification of holes on the reconstructed surface (Section B.2) and surface extraction in mesh generation (Section B.3).
B.2 Structured surface completion
In this section, we outline the detailed algorithms for our structured surface completion strategy. Algorithm 1 provides the pseudo-code for calibrating and identifying surface boundaries. Initially, we eliminate short line segments, which, despite their usefulness for tracking, are often considered noisy and less representative of the primary axes of the scene. By clustering the remaining longer segments, we identify the principal axes of the scene, aligning them with the three-dimensional coordinate system of the world through a calibration matrix. This calibration process ensures that the scene adheres to the Manhattan World assumption, where structured surfaces such as walls and floors are either parallel or perpendicular to the world coordinate axes.
Subsequently, we define the hypothesis surface, which acts as the ideal complete surface for interpolating holes. Specifically, we project 3D line features onto the plane $\bar{z}$, using the boundaries of these line features to define the hypothesis surface. Figure 7 illustrates the ablation of our calibration and surface identification method. Without calibration, we use the boundaries of $\mu_{\mathcal{G}}$, which are associated with structured surfaces, as the plane boundary and the mean value of $z\in\mu_{\mathcal{G}}$ for height. However, this approach results in misaligning the hypothesis surface with the actual floor. We also present a result using PCA on the centers of target Gaussians $\mu_{\mathcal{G}}$ to identify primary axes and calibrate the scene. Despite this effort, the predicted surface remains inaccurately positioned relative to the target surface, particularly showing disparities along the edges. And this disparity increases as the gaps on the surface become larger. By contrast, utilizing concrete 3D line features, which accurately represent scene boundaries and reflect primary axes, ensures precise identification of structured surfaces.
After calibration, we strategically interpolate new Gaussians to fill the holes and gaps in structured surfaces. The pseudo-code for our scene completion process is detailed in Algorithm 2. Holes on the hypothetical surface are specifically identified by the density function $\nu^{*}(p)$, as defined in Equation 13, where $\nu^{*}(p)$ falls below a threshold $\mathcal{T}_{density}$. To seamlessly fill these holes with appropriate texture, we train a PointNet++ model [34] that takes the Gaussian center $\mu{\mathcal{G}_{i}}$ as input and outputs texture colors. This approach is based on the understanding that surface textures in indoor scenes are regular and can be effectively predicted using spatial correlations. The training involves using Gaussians that are randomly sampled from the extracted incomplete surfaces, and we predict the texture color of interpolated Gaussians positioned at the holes. The scales of the generated Gaussians in the first two dimensions are determined based on the mean values of all existing Gaussians on the surface, while the third dimension is set to zero. This ensures the generated Gaussians are flat, facilitating the mesh extraction process.
We employed the PointNet++ model [34] for its proven efficiency in processing point clouds, capturing both local and global structural information. This model is ideally suited for the Manhattan World hypothesis, where the consistent flatness and regulation of surfaces simplify the task of texture interpolation. The straightforward yet effective architecture of the PointNet++ [34] allows it to discern detailed spatial patterns, seamlessly fits our framework to provide robust prediction without the computational overhead of more complex models.
1:Line segments $L\in\mathbb{R}^{3}$ and the reconstructed Gaussian scene $\mathcal{G}$
3:for each line segment $\ell$ in $L_{\rm filtered}$ do
4:     $\{direction\}\leftarrow(\ell.end-\ell.start).normalized()$triangleright Add direction vector
5:end for
6:$\vec{x},\vec{y},\vec{z}\leftarrow k\_means(\{direction\},k=3)$triangleright Cluster directions to find three main axes
7:$K\leftarrow\mathbb{I}\in\mathbb{R}^{4}$triangleright Define calibration matrix
@@ B.2 Structured surface completion
[…]
11:for each $\mathcal{G}_{i}$ in $\mathcal{G}$ do
12:     $\mu_{\mathcal{G}_{i}}\leftarrow K\cdot\mu_{\mathcal{G}_{i}}$triangleright Calibrate coordinates of Gaussian
13:     $\gamma_{\mathcal{G}_{i}}\leftarrow K\cdot\gamma_{\mathcal{G}_{i}}$triangleright Calibrate rotation of Gaussian
14:end for
15:for each line segment $\ell$ in $L$ do
16:     $\ell\leftarrow K\cdot\ell$triangleright Calibrate line segment
17:end for
$\mathcal{B}\leftarrow$ select maximum and minimum $x$ and $y$ for endpoints in $L$triangleright Define surface boundary
19:Output: boundary of surface $\mathcal{B}\in\mathbb{R}^{2}$
Algorithm 2 The pseudo-code for scene completion
3:for each surface $\mathcal{G}_{i}$ in $\mathcal{G}(\mathcal{M}_{tg})$ do
4:     $\mathcal{G}_{i}\leftarrow DBSCAN.fit(\mathcal{G}_{i})$triangleright use DBSCAN [13] to eliminate outliers
5:     $\bar{z}\leftarrow avg(z)$ for $z\in\mu_{\mathcal{G}_{i}}$
6:     $\Pi\leftarrow grid(\mathcal{B})$triangleright Create the 2D grid of hypothesis surface
7:     $\mathcal{D}_{i}\leftarrow$ compute the Gaussian density for each point of the grid in $\Pi$.
8:     if $\mathcal{D}_{i}$ < $\mathcal{T}_{density}$ then
9:         generates a new Gaussian onto the grid of low-density
10:         $\mu_{pred}\leftarrow$ records the center coordinates of the generated new Gaussian.
11:     end if
13:     for each batch $B_{i}$ in $B^{*}_{\mathcal{G}_{i}}$ do
15:     end for
16:     $\mathcal{G}_{\rm pred}\leftarrow$ predict the color using $\mathcal{F}$ using the center of interpolated Gaussian in $\mu_{pred}$
17:     $\mathcal{G}_{i}\leftarrow add~{}\mathcal{G}_{\rm pred}$
18:end for
19:$\mathcal{G}\leftarrow$ update the overall scene
20:Output: structure optimized $\mathcal{G}$
B.3 Mesh generation
We employed the surface extraction method proposed by [15] with an additional normal regularization term during the scene refinement. To compress the Gaussians to align closely with the surface, [15] forces the Gaussians to be flat by setting one of their scaling factors nearly to zero. Consequently, $\nu^{*}(p)$ can be further simplified to $\hat{\nu}^{*}(p)=\alpha_{\mathcal{G}^{*}}\mathrm{exp}\left(-{2s^{-2}_{\mathcal{G}^{*}}}\langle p-\mu_{\mathcal{G}^{*}},n_{\mathcal{G}^{*}}\rangle^{2}\right)$ where $s_{\mathcal{G}}$ represents the smallest scaling factor and $n_{\mathcal{G}}$ its corresponding normal. Here, $\alpha_{\mathcal{G}^{*}}=1$ is set to avoid semi-transparent Gaussians.
Moreover, to derive smooth structured surfaces, particularly in typically flat and textureless floor areas, we adopt the Manhattan World hypothesis, which assumes that planar floors are orthogonal to the vertical axis. We enforce this constraint by introducing a normal regularization term for the estimated SDF
In this context, the projected points $x_{\rm 2D}$, which pass through the camera ray and fall within the floor area, are encouraged to align their normal to the ideal vector $\hat{n}=\langle 0,0,1\rangle$. The set $\mathcal{P}_{\text{floor}}$ consists of image pixels identified as the floor using the semantic segmentation described in Section 3.3.2.
Figure 8 visualizes the outcomes of surface extraction with and without the normal regularization loss. It can be observed that incorporating this term enables a smoother reconstruction of the floor. In Figure 9, we compare the extracted mesh from our system to the NeRF-based approaches. Although the mesh generated by our method does not outperform recent NeRF-based methods [18, 36] in terms of quality, it still achieves good geometric accuracy and effectively produces mesh for Gaussian SLAM systems, which was not available in previous Gaussian SLAM methods.
Figure 8: The ablation of normal regularization for surface extraction. We compare the normal map of the extracted mesh, both without and with normal regularization, to the ground-truth mesh. The regularization refines the planar surfaces of the scene, i.e. floors, effectively.
Figure 9: Comparison of the mesh quality of our method with NeRF-based approaches on the Replica dataset [38]. Previous Gaussian-based methods cannot effectively generate mesh, and therefore, cannot be compared here.
Appendix C Ablation study
@@
In this section, we provide a comprehensive ablation analysis of each component’s effect and the hyperparameters of our model.
C.1 Ablation of hyperparameters
Our tracking system, which advances the foundational PLVS system [14], utilizes feature points and fused line segments for optimizing camera poses and performing bundle adjustments. The upper row of Figure 10 illustrates the effects of varying the number of points and line segments on the ATE loss [cm] and the frame rate of the tracking system. We identified an optimal region where having too few points and lines results in a lack of sufficient anchor features, whereas too many points and lines can lead to notable feature mismatches that reduce tracking accuracy. The tracking FPS reduces as a tradeoff with the increase in the number of features. Compared to tracking solely by point features, including line segment features noticeably decreases the frame rate; however, it still maintains a high rate over 30 FPS and does not become the speed bottleneck of the overall systems. The bottom row of Figure 10 illustrates the effects of default keyframe intervals and downsample ratio on the Gaussian map. During the mapping process, we uniformly downsample the point cloud generated from RGB-D input by a specific ratio to accelerate the system and conserve memory. We observed a marginal PSNR disparity when maintaining a relatively small downsample ratio.
@@ C.1 Ablation of hyperparameters
Figure 10: The ablation study examining the impact of the number of points, line segments, keyframe intervals, and downsample ratios in MG-SLAM on scene $0$_00 from the ScanNet dataset [6].
C.2 Ablation of tracking and mapping loss
MG-SLAM employs line segments for robust camera pose optimization and fine-grained map reconstruction. Specifically, fused line segments are utilized in the bundle adjustment of the tracking procedure, and a photometric loss related to line features is integrated into the map optimization. We use scene $0$_00 from the ScanNet dataset [6] to conduct the ablation study that evaluates the impact of each loss. This scene was chosen because it contains rich line features, which allow for a clear assessment of each component’s effectiveness. Table 5 presents the ATE [cm] and PSNR [dB] in relation to the use of each loss. We observed that integrating line features remarkably enhances tracking accuracy, and the fusion of line segments, which facilitates robust edge extraction, further improves the performance. For mapping, the photometric loss provides additional geometric constraints, thereby offering optimal reconstruction quality.
In Figure 11, we present a typical case in scene $0$_00 to demonstrate the effectiveness of each loss component. This case features a bike that appears multiple times along the camera trajectory. Compared to relying solely on point features, adding fused line segments significantly improves the quality of scene reconstruction. This enhancement mitigates scene drift on the bike and floor by providing more accurate camera pose estimation. Moreover, incorporating the photometric loss of line features in mapping results in more detailed object-level geometry and more precise reconstruction of line-like textures.
@@ C.2 Ablation of tracking and mapping loss
Table 5: Quantitative comparison of the ablation study on line losses, measuring ATE [cm] and PSNR [dB], conducted on scene $0$_00 of the ScanNet dataset [6]. The checkmark symbol indicates the employment of the method.
@@
Figure 11: The visualization of ablation analysis evaluating the effect of line segment losses in MG-SLAM. From left to right, the systems are displayed starting completely without line features, then adding line segments, integrating segment fusion, and finally incorporating line photometric loss in mapping. Key differences are highlighted with colored boxes.
Appendix D More experimental results
D.1 Training view evaluation on Replica-V1 dataset
Table 6 provides a detailed quantitative analysis of our training view rendering on the Replica-V1 dataset [38]. Our approach achieves state-of-the-art results, outperforming existing Gaussian-based methods by a remarkable margin across three average metrics.
Table 6: Quantitative comparison of our method and the baselines in training view rendering on the Replica-V1 dataset [38]. Our method demonstrates SOTA performances in most cases among three metrics.
[…]
D.2 Novel view synthesis on the large scale Replica Apartment dataset
We present additional experimental results on the Replica Apartment dataset [38], synthesized by [21]. This dataset poses an extremely challenging test case for visual SLAM systems due to its large scenes encompassing multiple rooms, each filled with numerous objects featuring complex geometries. Additionally, the camera trajectories include looping paths across multiple rooms, which introduces difficulties in maintaining consistent localization and mapping over extended periods.
We compare our method against SplaTAM [19] and MonoGS [28], which are Gaussian-based methods that have demonstrated impressive performance on the synthetic Replica-V1 dataset [38]. Figure 12, Figure 13, Figure 14, and Figure 15 visualize the mapping results, featuring the captured line segments and the final reconstruction outcomes on the left. On the right, comparisons with baseline methods and ground-truth are presented, with regions of interest highlighted with colored boxes.
Specifically, Figure 12 displays the reconstructed map of an extensive large scene containing two floors, eight rooms, and a staircase. SplaTAM [19] fails to track the camera trajectory, resulting in an entirely erroneously constructed second floor. MonoGS [28] establishes the whole map but struggles to capture the fine details of the environment and poses incorrect geometry of the objects and surfaces. In contrast, our method provides robust scene reconstruction with accurate geometry, which proves its effectiveness in extremely large indoor scenes.
Figure 13 shows that SplaTAM [19] struggles to accurately estimate the camera trajectory through two connected rooms, resulting in significant scene drift in the hallway. MonoGS [28] presents reliable pose estimation but faces challenges in reconstructing dense structural surfaces like walls, leaving gaps in these areas. In Figure 14, we observed that the map reconstructed by SplaTAM [19] includes many noisy Gaussians, likely because its abundant isotropic Gaussian representation cannot be efficiently optimized under limited camera angles. MonoGS [28] also encounters difficulties in reconstructing objects seen in the looping trajectory, such as the round table, due to inaccurate estimations of bundle adjustment across multiple rooms. Moreover, Figure 15 demonstrates a scene with a large camera rotation in the trajectory, under which MonoGS [28] produces incorrect geometry due to inaccurate camera tracking. Similarly, SplaTAM [19] displays noisy Gaussians in certain scenarios. Conversely, our method consistently delivers high-fidelity maps, accurately capturing scene geometry. This effectiveness can be attributed to the precise extraction and integration of line segments, which provides solid feature foundations for robust camera tracking and high-quality reconstruction in large scenes.
Figure 12: The novel view evaluation of the scene apartment 0 from the Replica Apartment dataset [38].
Figure 13: The novel view evaluation of the scene apartment 1 from the Replica Apartment dataset [38]. The top left shows the line segments extracted in 3D space. The bottom left illustrates the scene as reconstructed by MG-SLAM.
Figure 14: The novel view evaluation of the scene apartment 2 from the Replica Apartment dataset [38].
Figure 15: The novel view evaluation of the scene frl apartment 4 from the Replica Apartment dataset [38].
D.3 Evaluation of localization
In this section, we provide additional tracking evaluation of traditional SLAM systems and our method using the ScanNet dataset [6] and the Replica Apartment dataset [38].
On the ScanNet dataset [6], as shown in Table 7, our system shows optimal tracking performance over the ORB-SLAM systems [31, 2], which are known for their robustness in feature-based tracking and mapping, and PLVS [14], which forms the basis of our tracking systems.
On the Replica Apartment dataset [38], we compare MG-SLAM with both traditional and neural dense SLAM systems. We only evaluate ESLAM [18] as a NeRF baseline here since Point-SLAM [36], discussed in our main text, operates extremely slowly and struggles in large scenes. For Gaussian-based methods, our comparisons include SplaTAM [19] and MonoGS [28]. Table 8 shows that MG-SLAM achieves minimal ATE loss compared to traditional and neural dense baselines. In extremely large scenes, such as apartment 0, ORB-SLAM systems [31, 2] lose the camera trajectory on typical textureless areas such as the staircase, and require relocalization. In contrast, MG-SLAM, which utilizes fused line segments as robust anchor features, exhibits SOTA tracking performance. In addition, it is worth recalling that our method offers dense and high-fidelity map reconstruction which is not available in traditional SLAM systems.
Table 7: Additional quantitative comparison of our method and the traditional RGB-D SLAM methods in terms of ATE [cm] on the ScanNet dataset [6]. Our method demonstrates SOTA performances in all presented scenes. Comparison with other neural dense SLAM systems is outlined in the Table 2 of the main text.
  Methods Avg. $0$ 0059 0106 0169 0207
ORB-SLAM2 [31] 8.70 8.04 6.82 9.22 10.15 9.25
ORB-SLAM3 [2] 8.43 7.62 7.94 9.36 10.0 7.23
PLVS [14] 7.43 6.58 6.80 8.73 8.45 6.62
Ours 6.77 5.95 6.41 8.07 7.29 6.14
Table 8: Quantitative comparison of our method with both traditional and neural dense RGB-D SLAM systems in terms of ATE [cm] on the Replica Apartment dataset [38]. Our method shows SOTA performances, particularly notable improvements in the extensive two-floor apartment 0 scene. The underlined values indicate that relocalization was necessary due to loss of tracking.
Dense        Trad. Methods Avg. apart.0 apart.1 apart.2 frl_apart.0 frl_apart.4
ORB-SLAM2 [31] 3.78 12.83 6.82 4.70 1.85 1.76
ORB-SLAM3 [2] 3.45 9.45 6.54 4.17 1.63 1.48
PLVS [14] 2.94 7.06 4.78 4.04 1.45 1.39
ESLAM [18] 7.08 17.33 13.80 7.65 3.21 3.67
SplaTAM [19] 6.97 25.14 15.67 6.17 2.74 3.31
MonoGS [2] 6.55 20.69 9.51 7.57 2.98 6.14
Ours 2.08 5.03 3.78 2.75 0.92 0.88
D.4 Evaluation of surface completion on the ScanNet dataset
Figure 16 presents the comparison of MG-SLAM with Gaussian-based methods and our approach without surface optimization. The Gaussian-based methods exhibit noticeable gaps, especially in floor areas, unable to capture the full scene structure. In contrast, our method, when combined with a surface optimization strategy, effectively addresses these issues by seamlessly filling in gaps, thus delivering a more coherent and detailed scene reconstruction.
Figure 16: Comparison of novel-view synthesis of our methods and Gaussian-based approaches and without surface optimization. Gaps are accentuated using red boxes.
D.5 Frame rate and parameter size
Processing speed and memory efficiency are pivotal for SLAM systems to effectively operate in real-time and scale in real-world environments. Table 9 presents the comparative performance of our method against neural dense baseline approaches. Notably, NeRF-based Point-SLAM and Gaussian-based SplaTAM demonstrate limitations in achieving appropriate real-time processing capabilities, with both showing considerably low speeds in tracking and mapping alongside higher memory consumption. In contrast, MG-SLAM, implemented in a multi-processing system, exhibits exceptional superiority in tracking performance with the highest frame rate reaching 35 FPS. This improvement in tracking efficiency can be attributed to the incorporation of traditional feature-based optimization, which does not require iterative rendering and computing photometric loss for camera pose optimization, a common practice in neural dense systems, thereby significantly enhancing the tracking frame rate.
MG-SLAM also achieves high-speed and high-fidelity mapping, accelerated by a downsampling strategy that effectively reduces the number of Gaussians without compromising rendering performance. The reconstructed scenes with fewer Gaussians also ensure moderate memory consumption, making it suitable for deployment in large-scale scenarios.
Table 9: System comparison in terms of tracking, mapping, and rendering FPS, and memory usage between our method and the baselines on the Replica dataset [38]. The values represent the average outcomes across 8 scenes. SOTA performances are highlighted.
  Methods Track. FPS [f/s]$\uparrow$ Map. FPS [f/s]$\uparrow$ Render. FPS [f/s]$\uparrow$ SLAM FPS [f/s]$\uparrow$ Param. [mb]$\downarrow$
ESLAM 9.92 2.23 2.5 1.82 73.7
Point-SLAM 0.42 0.06 1.3 0.05 69.6
SplaTAM 0.35 0.22 122 0.14 277.1
MonoGS 2.57 3.85 769 1.80 33.4
Ours 35.32 3.96 324 3.12 78.5
Appendix E Limitation
MG-SLAM has certain limitations. The scene completion strategy, which relies on extracted line segments and the planar surface assumption, is only effective for large structured surfaces. It struggles to effectively interpolate unobserved geometry on complex objects within the scene. However, this problem has long been recognized as extremely challenging in the field of 3D reconstruction and SLAM systems. We believe that our efforts are making meaningful progress and pushing the boundaries further in this area.

### v3

@@ IV-D Evaluation on ScanNet Dataset
TABLE VII: Quantitative comparison of our method and the neural dense baselines in terms of ATE [cm] on the ScanNet dataset [27].
  Methods Avg. $0$ 0059 0106 0169 0207
NeRF Co-SLAM [3] 8.46 11.14 9.36 5.90 7.14 8.75
ESLAM [4] 7.68 8.47 8.70 7.58 7.45 6.20
Point-SLAM [6] 11.68 10.24 7.81 8.65 22.16 9.54
Gaussian SplaTAM [20] 12.04 12.83 10.10 17.72 12.08 7.46
MonoGS [21] 13.86 15.94 13.03 19.44 10.44 10.46
Photo-SLAM [24] 8.40 7.62 7.76 9.36 10.01 7.23
RTG-SLAM [39] 8.70 8.04 6.82 9.22 10.15 9.25
Ours 6.77 5.95 6.41 8.07 7.29 6.14
We provide quantitative assessments of reconstruction quality using the ScanNet dataset [27] in Table VI. Our approach delivers state-of-the-art results, outperforming other Gaussian-based methods by a notable 3dB in PSNR in real-world environments. The tracking evaluation results are shown in Table VII. Our method remarkably reduces the ATE RMSE (cm) error, achieving 20% improvements over baseline approaches.
In Figure 4, we visualize the novel-view synthesis results for MG-SLAM, comparing it with both NeRF-based and Gaussian-based SLAM systems. Our method exhibits robust and high-fidelity reconstruction capabilities. In comparison to Point-SLAM [6], our method offers complete scene reconstructions and finer texture details, benefiting from the use of Gaussians to handle complex geometries. Among Gaussian-based methods, SplaTAM [20] struggles in large scenes with complex camera loops due to its lack of bundle adjustment in the tracking system. Similarly, MonoGS [21] generally delivers reliable quality but struggles with object-level reconstruction drift. Photo-SLAM [24] achieves better reconstruction results compared to other baseline approaches but suffers from floaters and artifacts that compromise its geometric accuracy. In contrast, MG-SLAM excels with robust tracking and effective bundle adjustment incorporating line segments, enabling superior detailed reconstructions even surpassing the ground-truth mesh derived from offline methods.
IV-E Evaluation on Physical Platform
Figure 5: Qualitative comparison of our method and the baselines on the trajectory collected using our physical platform. The left side displays the line and point features extracted by our tracking system. The right side presents the reconstruction comparisons with Gaussian-based methods, with MG-SLAM showing more reliable reconstruction results in long-horizon and textureless indoor environments.
TABLE VIII: Quantitative comparison of our method and Gaussian-based approaches in terms of tracking and mapping performance on trajectory collected using our physical platform. The dash indicates system failures.
@@ IV-E Evaluation on Physical Platform
SplaTAM [20] - - - -
MonoGS [21] 27.04 18.28 0.751 0.677
Photo-SLAM [24] 18.29 22.92 0.798 0.620
Ours 7.72 25.47 0.831 0.575
To assess our system’s performance in real-world environments, we utilzie a private data trajectory gathered with our physical platform. The robot was deployed in the NTU building to navigate and collect data from a challenging long corridor, notable for its lack of texture and extensive length exceeding 200 meters. Ground-truth poses were computed by registering the LiDAR point cloud with the scanned point cloud generated by the Leica ScanStation. Figure 5 displays the comparisons of training view synthesis between our method and Gaussian-based systems, where our method demonstrates significant advantages; however, it still lacks some details compared to the ground-truth images due to the long-horizon trajectory. Table VIII provides quantitative results, showing that our method achieves more accurate tracking and mapping performance. The baseline approaches struggle with tracking in textureless environments, further affirming the effectiveness of our method in indoor scenes.
IV-F Evaluation of Scene Completion
The Gaussian SLAM faces limitations in interpolating the geometry of unseen views. This issue is especially evident in indoor scenes that feature complex layouts, where basic surfaces like floors are often obscured and poorly represented. Figure 6 qualitatively compared our method with recent Gaussian-based approaches in the Replica scenes [73] that presented uncovered geometry in camera trajectories. Utilizing the surface interpolation strategy based on the MW hypothesis [28], our method can accurately detect and proactively generate new Gaussians efficiently to fill gaps with certain textures, whereas Gaussian baseline methods exhibit substantial defects on structured surfaces.
Figure 6: Qualitative comparison on the novel-view synthesis of our method and the Gaussian SLAM baselines for hole fillings on the Replica dataset [73]. Our method demonstrates superior capability in interpolating and filling gaps on structured planar surfaces, such as floors and ceilings.
V Ablation Study
@@
In this section, we provide a comprehensive ablation analysis of the hyperparameters and each component’s effect.
@@
Figure 7: The ablation study examining the impact of the number of points, line segments, keyframe intervals, and downsample ratios in MG-SLAM on scene0000_00 from the ScanNet dataset [27].
V-A Ablation of Hyperparameters
Our tracking system, which advances the foundation of PLVS system [13], utilizes feature points and fused line segments for optimizing camera poses and performing bundle adjustments. The upper row of Figure 7 illustrates the effects of varying the number of points and line segments on the ATE loss [cm] and the frame rate of the tracking system. We identified an optimal region where having too few points and lines results in a lack of sufficient anchor features, whereas too many points and lines can lead to notable feature mismatches that reduce tracking accuracy. The tracking FPS reduces as a tradeoff with the increase in the number of features. Compared to tracking solely by point features, including line segment features noticeably decreases the frame rate; however, it still maintains a high rate over 30 FPS and does not become the speed bottleneck of the overall systems. The bottom row of Figure 7 illustrates the effects of default keyframe intervals and downsample ratio on the Gaussian map. During the mapping process, we uniformly downsample the point cloud generated from RGB-D input by a specific ratio to accelerate the system and conserve memory. We observed a marginal PSNR disparity when maintaining a relatively small downsample ratio.
V-B Ablation of Line Segment Extraction
Figure 8: The ablation study of the line segment extraction outcomes for our method on the scene frl_apartment_4 from the Replica Apartment dataset [73]. By incorporating backprojection, filtering, and segment fusion, our system generates rich and robust line segment features for subsequent tracking and mapping procedures.
The extracted line segments play a crucial role in serving as robust feature foundations in the subsequent optimization processes. Figure 8 illustrates the ablation results for our feature extraction approach for these 3D line segments, demonstrating the efficacy of segment backprojection, line fusion, and filtering strategies in providing clear and accurate line features.
Figure 9: The visualization of the ablation study in scene0000_00 from the ScanNet dataset [27]. From left to right, the results are displayed starting completely without line features, adding line segments, integrating segment fusion, and finally incorporating line photometric loss in mapping. Key differences are highlighted with colored boxes.
V-C Ablation of Tracking and Mapping Loss
@@
TABLE IX: Quantitative comparison of the ablation study on line losses, measuring ATE [cm] and PSNR [dB], conducted on scene0000_00 of the ScanNet dataset [27]. The checkmark symbol indicates the employment of the method.
@@
MG-SLAM employs line segments for robust camera pose optimization and fine-grained map reconstruction. Specifically, fused line segments are utilized in the bundle adjustment of the tracking procedure, and a photometric loss related to line features is integrated into the map optimization. We use scene0000_00 from the ScanNet dataset [27] to conduct the ablation study that evaluates the impact of each loss. This scene was chosen because it contains rich line features, which allow for a clear assessment of each component’s effectiveness. Table IX presents the ATE [cm] and PSNR [dB] in relation to the use of each loss. We observed that integrating line features remarkably enhances tracking accuracy, and the fusion of line segments, which facilitates robust edge extraction, further improves the performance. For mapping, the photometric loss provides additional geometric constraints, thereby offering optimal reconstruction quality.
In Figure 9, we present a typical case in scene0000_00 to demonstrate the effectiveness of each loss component. This case features a bike that appears multiple times along the camera trajectory. Compared to relying solely on point features, adding fused line segments significantly improves the quality of map reconstruction. This enhancement mitigates scene drift by providing more accurate camera pose estimation. Moreover, incorporating the photometric loss of line features in mapping results in more detailed object-level geometry and more precise reconstruction of line-like textures.
V-D Ablation of Surface Extraction
We employ hypothesis surfaces regularized by line segments to pinpoint potential gaps on structured surfaces, as described in Section III-D. These identified hypothesis surfaces are designed to precisely mirror the positions of the planar surfaces with minimal redundancy. To delineate these surfaces, we initially adjust the reconstructed map to adhere to the orthogonality assumption of MW. Figure 10 displays the hypothesized rectangular planes produced without calibration, with PCA calibration based on the centers of Gaussian primitives, and with calibration using features from extracted segments. It is evident that our segment-based method efficiently captures the floor surface with reduced redundancy.
Figure 11: The ablation of normal regularization for surface extraction on the Replica dataset [73]. We compare the normal map of the extracted mesh, both without and with normal regularization, to the ground-truth mesh. The regularization refines the planar surfaces of the scene, i.e. floors, effectively.
V-E Ablation of Mesh Generation
Following map reconstruction and surface interpolation, we extract meshes from the Gaussian map, as outlined in Section III-E. Figure 11 illustrates the results of mesh extraction with and without the proposed normal regularization loss. We observe that incorperating this regularization term results in smoother mesh extraction on the structured surface such as floors compared to the original method introduced by [31].
VI Conclusion
In this study, we present MG-SLAM, a Gaussian-based SLAM method based on the MW hypothesis. MG-SLAM employs fused line segments and point features for robust pose estimation and map refinement. Furthermore, by leveraging the line segments and planar surface assumption, we efficiently generate new Gaussian primitives in gaps on structural surfaces caused by obstructions or unseen geometry. Extensive experiments demonstrate that our method delivers state-of-the-art tracking and mapping performance, while also maintaining real-time processing speed.
VII Limitations
MW assumes surface planes generally align with the three orthogonal directions defined by the Cartesian coordinate system. This presents challenges in refining planes or layouts that are slanted or not strictly orthogonal. To better accommodate diverse outdoor urban environments, future research could incorporate multiple horizontal dominant directions, as seen in Atlanta World [78], include additional sloping directions like those in Hong Kong World [79], or adopt uniform inclination angles for the dominant directions proposed in San Francisco World [80]. Additionally, since the optimization of our dense map primarily relies on line features and large planar surfaces, it struggles with refining piece-wise or fine-grained structural planes, such as walls and windows. One potential approach is to leverage the coplanarity of surfaces for structural regularization [61, 81], although this proves particularly challenging in indoor environments with complex layouts where identifying coplanar surfaces is difficult.